# Modular analogs of character formulas and multiplicities of induced representations

### v1

@@ A.4. Quaternion division algebras
Lemma A.17 ([Vig89, Proposition 9(a), p. 260], cf. [Tok15, Proposition 2.5]).
@@
Proof of Corollary A.16.
We will show that, for every character $\epsilon$ of $D^{\times}$, the space $\mathrm{Hom}_{D^{\times}}\mathopen{}\left(\pi_{1}\otimes\pi_{2}\otimes\pi_{3},\epsilon\right)\mathclose{}$ has dimension at most one. Since, as discussed above, each $\pi_{i}$ has a twist that is inflated from $G=D^{\times}/\varpi_{F}^{\mathbb{Z}}U_{D}^{1}=C\rtimes\mathopen{}\left\langle\sigma\right\rangle\mathclose{}$, and since this proposition is invariant under twisting, we may assume that each $\pi_{i}$ is inflated from $G$. Moreover, if at least two of $\pi_{1}$, $\pi_{2}$ and $\pi_{3}$ are characters, then $\pi_{1}\otimes\pi_{2}\otimes\pi_{3}$ is irreducible and there is nothing to show, so we assume, without loss of generality, that both $\pi_{1}$ and $\pi_{2}$ have dimension greater than one. We are now working in the semisimple category of $k[G]$-modules (since $p\nmid\#G)$, and we have to show that the multiplicity of the representation $\pi_{3}^{\vee}\otimes\epsilon$ in $\pi_{1}\otimes\pi_{2}$ is at most one, where $\pi_{3}^{\vee}$ is the contragredient of $\pi_{3}$. In fact, $\pi_{1}\otimes\pi_{2}$ is multiplicity free, as we now show.
By Lemma A.17, there are characters $\psi_{1},\psi_{2}:E^{\times}\to k^{\times}$ with $\psi_{i}\neq\psi_{i}^{\sigma}$ such that $\pi_{i}=\mathrm{Ind}_{E^{\times}U_{D}^{1}}^{D^{\times}}(\psi_{i}^{\dagger})$. Identifying each $\pi_{i}$ with the associated representation of $G$, this is simply $\pi_{i}=\mathrm{Ind}_{C}^{G}\psi_{i}$. Then
@@
by the push-pull formula and Mackey’s formula. Since $\psi_{i}\neq\psi_{i}^{\sigma}$, it follows that $\psi_{1}\psi_{2}\neq\psi_{1}\psi_{2}^{\sigma}$ and $\psi_{1}^{\sigma}\psi_{2}^{\sigma}\neq\psi_{1}\psi_{2}^{\sigma}$, so (using Lemma A.17 again) this is a direct sum of two distinct two-dimensional representations that do not share a common subrepresentation. In particular, $\pi_{1}\otimes\pi_{2}$ is multiplicity free.

### v2

@@ 4. Quaternion division algebras
Lemma 4.2 ([Vig89, Proposition 9(a), p. 260], cf. [Tok15, Proposition 2.5]).
@@
Proof of Corollary 4.1.
We will prove the following stronger statement: for every character $\epsilon$ of $D^{\times}$, the space $\mathrm{Hom}_{D^{\times}}\mathopen{}\left(\pi_{1}\otimes\pi_{2}\otimes\pi_{3},\epsilon\right)\mathclose{}$ has dimension at most one. Since, as discussed above, each $\pi_{i}$ has a twist that is inflated from $G=D^{\times}/\varpi_{F}^{\mathbb{Z}}U_{D}^{1}=C\rtimes\mathopen{}\left\langle\sigma\right\rangle\mathclose{}$, and since this statement is invariant under twisting, we may assume that each $\pi_{i}$ is inflated from $G$. Moreover, if at least two of $\pi_{1}$, $\pi_{2}$ and $\pi_{3}$ are characters, then $\pi_{1}\otimes\pi_{2}\otimes\pi_{3}$ is irreducible and there is nothing to show, so we assume, without loss of generality, that both $\pi_{1}$ and $\pi_{2}$ have dimension greater than one. We are now working in the semisimple category of $k[G]$-modules (since $p\nmid\#G)$, and we have to show that the multiplicity of the representation $\pi_{3}^{\vee}\otimes\epsilon$ in $\pi_{1}\otimes\pi_{2}$ is at most one, where $\pi_{3}^{\vee}$ is the contragredient of $\pi_{3}$. In fact, $\pi_{1}\otimes\pi_{2}$ is multiplicity-free, as we now show.
By Lemma 4.2, there are characters $\psi_{1},\psi_{2}:E^{\times}\to k^{\times}$ with $\psi_{i}\neq\psi_{i}^{\sigma}$ such that $\pi_{i}=\mathrm{Ind}_{E^{\times}U_{D}^{1}}^{D^{\times}}(\psi_{i}^{\dagger})$. Identifying each $\pi_{i}$ with the associated representation of $G$, this is simply $\pi_{i}=\mathrm{Ind}_{C}^{G}(\psi_{i})$. Then
@@
by the push-pull formula and Mackey’s formula. Since $\psi_{i}\neq\psi_{i}^{\sigma}$, it follows that $\psi_{1}\psi_{2}\neq\psi_{1}\psi_{2}^{\sigma}$ and $\psi_{1}^{\sigma}\psi_{2}^{\sigma}\neq\psi_{1}\psi_{2}^{\sigma}$, so (using Lemma 4.2 again) this is a direct sum of two distinct two-dimensional representations that do not share a common subrepresentation. In particular, $\pi_{1}\otimes\pi_{2}$ is multiplicity-free. ∎